You can stuff envelopes twice as fast as your friend. Together, you can stuff 6750 envelopes in 4.5 hours. How long would it take each of you working alone to complete the job?
Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)
Answer by ikleyn(53751)

.
You can stuff envelopes twice as fast as your friend. Together, you can stuff 6750 envelopes in 4.5 hours.
How long would it take each of you working alone to complete the job?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


            It is so simple that can be easy solved mentally.


From the problem, you and your friend work as productive as 2+1 = 3 copies of your friend.


So, 3 persons, working as productive as your friend, will complete the job in 4.5 hours.


It means that your friend, working alone, will complete the job in 3*4.5 = 13.5 hours.


Hence, you, working alone, will complete the job in half of 13.5 hours, i.e. in

    13.5 / 2 = 6.75 hours = 6 hours and 45 minutes.

Solved.

The answer in the post by @lynnlo is incorrect.